OpenStudy (anonymous):

find polynomial function that has given zeros: 0,-3

OpenStudy (jdoe0001):

\(\bf x=\{ 0,-3\}\to \begin{cases} x=0\to {\color{purple}{ (x)}}=0\\ \bf x=-3\to x+3=0\to {\color{purple}{ (x+3)}}=0 \end{cases} \\ \quad \\ {\color{purple}{ (x)}}{\color{purple}{ (x+3)}}=0\implies (x)(x+3)=\textit{original polynomial}\)